OpenStudy (anonymous):

what is 66 2/3% in its simplest form as a fraction

OpenStudy (anonymous):

Basically all you do is multiply 66 and three to get 198. Then you add 2 to 198 to get 200. Put the 200 over the denominator and the improper fraction of 66 2/3 is 200/3. 66x3=198 198+2=200 66 2/3= 200/3 To put an improper fraction in simplest form, first figure out what is the biggest number that both the numerator and denominator can go into without leaving a remainder. For example, take 99/6. Three is the biggest number that goes into both numbers so divide 99 and 6 by three to get 33/2. In this case, 200/3 is in its simplest form.

OpenStudy (mathstudent55):

\(66 ~2/3\% = (66 + \dfrac{2}{3}) \% = \left(\dfrac{3}{3} \times \dfrac{66}{1} + \dfrac{2}{3} \right) \% = \left(\dfrac{ 198}{3} + \dfrac{2}{3}\right) \% = \dfrac{200}{3} \times \dfrac{1}{100} = \dfrac{2}{3}\)

OpenStudy (anonymous):

thankz